OpenStudy (anonymous):

I have a problem where I must find the Cartesian equation for: x=t^2 y=4+3t into x=Ay^2+By+C I have found C, but need help in finding the values for A and B

OpenStudy (anonymous):

it seems you are either missing some info here or im a bit more rusty on geometry than i thought, how exactly did you find C?

OpenStudy (anonymous):

I came up with t^2=A(4+3t)^2+B(4+3t)+C

OpenStudy (anonymous):

but i'm still kind of confused as to how I even got it ha

OpenStudy (anonymous):

it's the correct problem i checked

OpenStudy (anonymous):

yeah, you have 3 variables there and no way to eliminate any of them, so unless your proff gave you some format to put it into and then from that you know the points your trying to find?

OpenStudy (anonymous):

i mean the way to eliminate them is to use the t value and then deal with that bottom equation

OpenStudy (anonymous):

so lik eeither put everything in terms of x or y

OpenStudy (anonymous):

right but you can only ever solve the variables in terms of the other variables, meaning you can find A as a function of B and C, B(A,C) and C(A,B)

OpenStudy (anonymous):

yes i know that's why i need to find a way out by plugging in constants on both sides of the equation that make A or B go to zero in order to solve for the variables

OpenStudy (anonymous):

ahh, right. and then you just do the same thing to find A and B i think, set the ones your not finding to 0

OpenStudy (anonymous):

see i don't know how to do the rest i'm really cracked out and can't thinka nymore ha

OpenStudy (anonymous):

honestly you should confinr that with someone else too though, its been a while since iv done a problem like this

OpenStudy (anonymous):

if your finding A you just set B and C to 0 because they are intercepts, right?

OpenStudy (anonymous):

well i'll know if it's right because i have to type my answers into a computer system that will tell me if it's correct or not

OpenStudy (anonymous):

unlimited trials

OpenStudy (anonymous):

so you have a polynomial function of the 2nd degree which is a quadradic

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

question where do u go to school?

OpenStudy (anonymous):

this is a trickey little bastard, you have to change the form of the original equation. see If b = 0, the quadratic function has the form f(x) = ax2 + c. so what i think you have to do, is noting those derived equations and plugging in 0 to the other variables you get some new equations which then allow you to solve for the others by plugging them into eachother. lemme keep reading

OpenStudy (anonymous):

ughhh it sucks!! i haven't slept in two days fml

OpenStudy (anonymous):

lol its been too long, i dont understand what that website is saying...

OpenStudy (anonymous):

ok let me think about this for a minute and see if i can figure it out

OpenStudy (anonymous):

do you work for this website or something or are you a student ? i don't really understand this website ha

OpenStudy (anonymous):

nah im just a student

OpenStudy (anonymous):

oh where?

OpenStudy (anonymous):

Rochester Institute of Technology

OpenStudy (anonymous):

small fluttering world

OpenStudy (anonymous):

i go to u of r

OpenStudy (anonymous):

no way hahaha thats crazy. i was about to say dont even lie to me that you go to rit, your sn has "girl" in it

OpenStudy (anonymous):

haha yea i don't go there a lot, ha or ever for that matter, but i have some friends from high school there

OpenStudy (anonymous):

oh your from around here, thats cool. what year are you?

OpenStudy (anonymous):

i'm a junior

OpenStudy (anonymous):

lol im a 3rd year too. i know a couple kids who went to school around here but i dont know what schools or anything cause im from boston

OpenStudy (anonymous):

bosston, i lived there this summer ha, well i lived in watertown

OpenStudy (anonymous):

literallrally everyone from u of r is from boston

OpenStudy (anonymous):

i say boston its really the southern NH border but i'v driven through watertown more than once. and yeah it seems like that here too, i know tons of people from the boston area

OpenStudy (anonymous):

im still in awe that at 2am out of 23 online people i found someone thats currently within a 10 mile radius of me

OpenStudy (anonymous):

i am too, it's just too weird

OpenStudy (anonymous):

what do you major in ?

OpenStudy (anonymous):

computer engineering technology, you?

OpenStudy (anonymous):

ha sounds intense, i'm doing business econ and psych

OpenStudy (anonymous):

haha im concentrating in psych, too weird

OpenStudy (anonymous):

haha oh the weird realms of the internet

OpenStudy (anonymous):

its got to be less coincidental than were making it out to be. were in a pretty tech-savvy part of the english speaking world. but still thats the whole northern east coast...

OpenStudy (anonymous):

how did you find this site?

OpenStudy (anonymous):

haha idk i was just wearching for help things and it came up

OpenStudy (anonymous):

why do you help random people with math?

OpenStudy (anonymous):

honestly i dont really know

OpenStudy (anonymous):

i kinda stumbled here like 2 days ago and got a kick out of helping people out with their math, most of the stuff on here is low level calc or easier so its not so bad

OpenStudy (anonymous):

guess i got sick of facebook and this was another way to procrastinate?

OpenStudy (anonymous):

i guess that's not a bad way to procrastinate, bettering the world's math skills

OpenStudy (anonymous):

haha, yeah doing other people's homework is fun but doing my own is agonizing, go figure

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

what class is your parabola thinger from? i dont remember having to do this stuff

OpenStudy (anonymous):

third semester of calc

OpenStudy (anonymous):

shows how much i learned in calc...

OpenStudy (anonymous):

but it's something i learned a long time ago, i just don't remember it

OpenStudy (anonymous):

don't worry though, i have the rochester minions working on it

OpenStudy (anonymous):

the rochester minions eigh?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you should prolly let me know how to go about contacting these minions. could be helpful for diff eq

OpenStudy (anonymous):

haha, you literally just walk around u of r campus, they're around, sometimes you have to look in the stacks or the depths of the librar but there's always a guarantee you'll find one

OpenStudy (anonymous):

haha, even at 2 am. that seems sketchy, especially the library part, whats a library?

OpenStudy (anonymous):

the minions are having trouble with it too, darn

OpenStudy (anonymous):

tell your professor that its impossible. ill see if my roomate knows how to do it

OpenStudy (anonymous):

ok it's just it's due tomorrow morning at 6am ick

OpenStudy (anonymous):

haha people keep looking at this question and then leaving

OpenStudy (anonymous):

theres 3 rit kids in here working on your silly parabola

OpenStudy (anonymous):

you said you have infinite number of trys?

OpenStudy (anonymous):

\[y=4+3\sqrt{x}\]

OpenStudy (anonymous):

give that a go

OpenStudy (anonymous):

i'll try it

OpenStudy (anonymous):

idk how long this website will wokrk though, its slowing my comp down a lot

OpenStudy (anonymous):

i got A it's 1/9, idk if that will help find B

OpenStudy (anonymous):

lol i have no idea how you got that or what it means. anyways im going to go sleep or eat or something, facebook me (Tim Heebs) you should come to some parties and bring friends. god knows RIT needs help with the ratio at parties haha

OpenStudy (anonymous):

haha yea, your name already showed up when i was worried about the problem, and yea us girls here sometimes want to go to some RIT parties, that would be cool

OpenStudy (anonymous):

sounds good, ttyl then. good luck with your calc